OpenStudy (anonymous):

A rectangle's width is 2ft less than its length. It's perimeter is 48 ft. Find the length of the rectangle. I got 12ft. Is this correct? Please and Thank you.

OpenStudy (goformit100):

See http://openstudy.com/study#/updates/5109ea33e4b0d9aa3c46552d

OpenStudy (zzr0ck3r):

x+ x + y + y = 48 -> 2x+2y = 48 where x = width and y = length x - 2 = y 2x + 2(x-2) = 48 4x - 4 = 48 -> x = 13 so y = 11 is what I get

OpenStudy (zzr0ck3r):

@kitsune0724 you see?

OpenStudy (anonymous):

Thank you. It's what I got after redoing the problem.